OpenStudy (anonymous):

Write a rule for the function represented in the table. x y 0 9 1 28 2 47 3 66

OpenStudy (anonymous):

Help? I tried and tried

OpenStudy (anonymous):

is y = 19 +9x

OpenStudy (anonymous):

my choices are: y = 9 + 19x y = 18 + 10x y = 19 + 9x y = 10 + 18x

OpenStudy (anonymous):

I think it is y = 19 +9x

OpenStudy (the_fizicx99):

M = y2 -y1 / x2 - x1 M = 47 - 28 / 2 - 1 M = 19/1 = 19 y-y1=m(x-x1) y - 28 = 19(x - 1) y - 28 = 19x - 19 +28 +28 y = 19x + 9 <- function rule Check: change y for f(x) and plug in 2 for x, you should get 47, if not then it doesn't work. f(x) = 19x + 9 f(2) = 19(2) + 9 --> 38 + 9 = 47 \(\ \color{green}{\checkmark}\) this rule works

OpenStudy (anonymous):

Wow! Thanks. Detailed. Ok going to read it just sec.

OpenStudy (the_fizicx99):

Close, it's y = 19x + 9 :P

OpenStudy (anonymous):

Glad I was at least close! Thanks! Can I ask one more ? I will close this and then ask on another so you can get more points.

OpenStudy (the_fizicx99):

Points don't really matter to be honest, it's only virtual! :P sure ask :>

OpenStudy (anonymous):

Ok thanks. Just sec